OpenStudy (anonymous):

22. A basketball is shot at a 70° angle with the horizontal with an initial velocity of 10 meters per second. Find the component form of the initial velocity. A. v ~ (9.40,3.42) B. v ~ (-3.42,9.40) C. v ~ (3.42,-9.40) D. v ~ (3.42, 9.40) E. v ~ (-9.40,-3.42) F. v ~ (9.40, -3.42)

OpenStudy (anonymous):

 component form requires the horizontal (x) and vertical (y) velocities how do you think you can obtain those (hint: the relationship can come form the picture)

OpenStudy (anonymous):

ok that didnt help me at all

OpenStudy (anonymous):

......

OpenStudy (anonymous):

i'm guessing this is for a trig class? you should have gone over how to relate the sides of a triangle to the hypotenuse. SOH CAH TOA ring a bell?

OpenStudy (anonymous):

ya it rings a bell

OpenStudy (anonymous):

so SOH CAH TOA is a memory aid which helps you remember how the sides of a right angle triangle are related through trig functions: example, consider SOH: \[\large \sin\theta = \frac{opposite}{hypotenuse}\] meaning that if you have an angle (in your case, 70 degrees) than the ratio of the side opposite the angle and the hypotenuse is given by sin(70)

OpenStudy (anonymous):

look at the equation, you'll see that there are 3 variables (opposite length, hypotenuse length, and the angle). If you are ever given two of the three you can solve the third by using this trig equation

OpenStudy (anonymous):

from the drawing you can see that you are given the hypotenuse (10m/s) and the angle (70 degrees. solving this will give you the opposite side which, by looking at the drawing, is the same as the vertical component of velocity: \[\large \sin(70^o)= \frac{opposite}{10}\] try solving that and see what you get.

OpenStudy (anonymous):

i think its D

OpenStudy (anonymous):

are you guessing or did you solve the triangle using trig?

OpenStudy (anonymous):

guessing, I dont understand but I gotta get this done tonight

OpenStudy (anonymous):

what part don't you understand?